OpenStudy (anonymous):

PLEASE HELP WILL GIVE MEDAL!! :) Find the vertices and foci of the hyperbola with equation ((x+1)^2)/16 - (y+5)^2/9 =1

OpenStudy (anonymous):

\[a^2=16\] a=4 \[b^2= 9\] b=3 the center is (1,5) (I got the 1 from the x+1 and the 5 from y+5) the vertices are 4 units from the center on either side of the x axis so that means it would be (5,5) and (-3,5) the foci are found with the equation \[c^2=a^2+b^2\] so C^2= 16 + 9 or C^2= 25 or c=5 so that means the foci are 5 units from the center (6,5) and (-5,5)

OpenStudy (anonymous):

did that help?

OpenStudy (anonymous):

http://www.purplemath.com/modules/hyperbola2.htm this might help